OpenStudy (anonymous):

A body of mass 2kg is flying horizontally with a velocity of 72km/hr at a height of 10m. Compute the total energy.

OpenStudy (anonymous):

hey u know the formula of total energy

OpenStudy (anonymous):

nop sry......

OpenStudy (anonymous):

ok total energy=kinetic +potential energy

OpenStudy (anonymous):

k

OpenStudy (anonymous):

total energy =1/2(mv^2) +mgh using this formula

OpenStudy (anonymous):

is 216 J the answer ??????

OpenStudy (anonymous):

*596J nt 216

OpenStudy (anonymous):

can u calculate

OpenStudy (anonymous):

i did...... jst to knw if it is right ??/

OpenStudy (anonymous):

TE = 1/2 mv^2 + mgh =1/2 x 2 x (20)^2 + 2 x 9.8 x 10 =400+196 =596J am i right????????????? O.o

OpenStudy (anonymous):

u there???????

OpenStudy (anonymous):

v=72 use

OpenStudy (anonymous):

but it is 72 km/hr......shoudn we convert it to metre per second (m/s)

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

any question

OpenStudy (anonymous):

so its correct right????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k hnx 4 ur help.... ^_^